# On Evaluating The Performance of Watermarked Machine-Generated Texts Under Adversarial Attacks

### v1

@@ Abstract
In this paper, to fill this gap, we first systematically comb the mainstream watermarking schemes and removal attacks on machine-generated texts, and then we categorize them into pre-text (before text generation) and post-text (after text generation) classes so that we can conduct diversified analyses. In our experiments, we evaluate eight watermarks (five pre-text, three post-text) and twelve attacks (two pre-text, ten post-text) across 87 scenarios. Evaluation results indicate that (1) KGW and Exponential watermarks offer high text quality and watermark retention but remain vulnerable to most attacks; (2) Post-text attacks are found to be more efficient and practical than pre-text attacks; (3) Pre-text watermarks are generally more imperceptible, as they do not alter text fluency, unlike post-text watermarks; (4) Additionally, combined attack methods can significantly increase effectiveness, highlighting the need for more robust watermarking solutions. Our study underscores the vulnerabilities of current techniques and the necessity for developing more resilient schemes.
@@ Introduction
Our work. In this paper, we fill the gap in the literature by systematically categorizing watermarks and attacks into pre-text and post-text classes, where the former involves watermark injection or disruption before text generation, and the latter does so after text generation. We consider eight watermarks (five pre-text and three post-text) and twelve attacks (two pre-text and ten post-text), resulting in a total of 87 possible scenarios (see Table 3). Our evaluation shows that KGW [29] and Exponential [1] are the two best watermarks by providing good text quality and relatively high watermark rates after different attacks compared to other watermarks. For instance, KGW and Exponential achieve over 0.5 robustness score while Unigram [79] and Linguistic [77] watermarks only reach around 0.3 robustness score (see Table 5 for more details). However, they are still not robust against most of the attacks. For example, KGW only reaches a 0.0349 watermark rate under Paraphrase attacks. This suggests that existing watermarks are still vulnerable to various attacks. Regarding efficiency, both KGW and Exponential are relatively efficient in watermark injection and detection. In terms of attacks, we find that post-text attacks are generally more efficient than pre-text attacks since they do not require modification of the model’s weights, making them more practical threats to current watermarks. Concerning imperceptibility, pre-text watermarks usually perform better as they are hidden in the token distributions and do not affect text fluency, whereas post-text watermarks involve adding, removing, or replacing tokens, making them easier to detect.
@@
  • We perform the most comprehensive (thus far) evaluation on the performance of watermarked machine-generated text under adversarial attacks.
@@
Following [53], we model the generation process of LLMs into two steps: probability prediction and token selection. First, given a sequence of tokens $x_{1:n}$ (e.g., the input prompt) where $x_{i}\in\{1,...,|V|\}$ is a token and $|V|$ is the size of vocabulary $V$, LLM calculates the conditional probability of the next token as $Pr(x_{n+1}|x_{1:n})$. Then, LLM uses a specific sampling strategy to select the word from the vocabulary as the final token, e.g., Greedy Sampling confirms next token by $x_{n+1}=\arg\max_{x}Pr(x|x_{1:n})$. Naturally, LLM generates a $l$ length text through computing $Pr(x_{1:n+l})=\prod_{i=1}^{n+l}Pr(x_{i}|x_{1:i-1})$.
@@
Black-box & White-box. A watermarking scheme is classified as white-box if it relies on the model’s internal parameters during watermark injection. Conversely, a scheme is black-box if it does not access model parameters. Notably, schemes accessing the last hidden state of an LLM are also considered as black-box.
@@
First of all, we clarify that we focus on the watermarking schemes which are based on the generated texts in this paper. Compared with the watermarking schemes which need to embed the watermarks by updating the model parameters [56, 36], these text-based schemes can be adapted to any causal language models. Meanwhile, we consider two types of watermarks, i.e., pre-text watermarks and post-text watermarks, where the former injects the watermark before or during text generation and the later injects the watermark after the text generation.
The methods of pre-text watermark can be divided into two categories based on the process used for generating text with an LLM [41]. For instance, this includes modifying the logits [29, 79, 40, 35, 45, 24, 43] and modifying the token sampling strategy during the inference phase [34, 1, 9]. The methods of post-text watermark can be divided into four categories based on the granularity of text modification [41], which are format-based watermark [58, 64], lexcial-based watermark [77, 51], synatic-based watermark [70, 48], and generation-based watermark [78, 2]. To comprehensively discuss the robustness of various types of watermarking schemes, we discuss eight LLM watermarking schemes in this paper, which are summarized in Table 1.
@@
Pre-text watermarks can be categorized into token sampling-based and logits modification-based watermarks where the former modifies the token sampling strategy and the later modifies logits during the inference phase.
@@
  • Convert bit-string (Convert) [9] aims to develop undetectable watermarks. This watermark can be viewed as the hash operation on the generated text. Specifically, they first encode the generated text into a string with binary-bits. Then, they use a {0,1} hash function $h$ and sample tokens $x_{j}$ with preference for those satisfying $h(x_{j})=1$. More tokens should hash to 1 than to 0 in watermarked text. The probability of a token being encoded to 1 is defined as $p_{j}(1)$. The watermarked model will output $x_{j}=1$ if $u_{j}\leq p_{j}(1)$ and $x_{j}=0$ otherwise. Here $u_{j}$ is a real number sampled from [0,1]. The probability that the watermarked model output $x_{j}=1$ is $p_{j}(1)$. For detection, they will compute a score $s(x_{j},u_{j})=\ln\frac{1}{u_{j}}$ if $x_{j}=1$ and $s(x_{j},u_{j})=\ln\frac{1}{1-u_{j}}$ otherwise. Given a string $(x_{1},...,x_{L})$, the sum scores of detection is defined as $score=\sum_{j=1}^{L}s(x_{j},u_{j})$. The score of the watermarked text should be significantly higher than the score of the normal text.
@@
  • Inverse token sampling (Inverse) [34] is similar to exponential. This strategy calculate the watermark scores from the next position of the watermark key $\xi$, $\xi=\xi^{(1)},\xi^{(2)},...,\xi^{(m)}$ where $\xi^{(i)}\in[0,1]^{\lvert\mathcal{V}\rvert}$ and m should exceed the max length of generated text, $\mathcal{V}$ is the vocabulary of tokens. Finally, the watermark strategy is defined as:
@@
  • KGW. Kirchenbauer et al. [29] propose a watermarking scheme without retraining the model. Specifically, during the inference process, a secret key can be calculated by the previous token and we can use the secret key to partition the vocabulary into a green list and a red list. Then, we add a bias to each green list logit before the Softmax function to calculate the probability distribution. We can detect the watermark by making a statistical test to the proportion of tokens in the text that belongs to green list. Finally, they statistical measure the z-score which is defined as:
@@
    where $|s|_{G}$ presents the number of green tokens, $L$ presents the total number of tokens in the text and $\gamma$ is a configuration parameter that presents the proportion of green list tokens in the whole vocabulary. We consider the text is watermarked if its z-score exceeds the pre-defind threshold.
@@
Post-text watermarks constitute a category of schemes wherein watermarks are appended to existing texts. The predominant approach for this type of watermark involves the modification of the existing text to incorporate the watermarking information. We implement three watermark schemes belonging to this category in this paper, and we divided the three watermarks into two categories which are Format-based watermark and Lexical-based watermark according to the granularity of modifications.
Format-based watermark. The Format-based watermark derives from a image watermark technology [5]. It does not modify the content of text but changes the format of text that are imperceptible for human, thereby implement the watermark injection.
  • WHITEMARK [64] exploits the fact that the Unicode of text has several codepoints for whitespace. They can replaces the origin whitespace (e.g. U+0020) to another new whitespace (e.g. U+2004) without degrading the quality of text. The watermark can be detected by calculating the probability of new whitespace.
@@
Lexical-based watermark. This catergory of watermarks apply word-level modifications by replacing the random sampling words with other candidate synonyms without affecting the structure of the whole sentence. This category of watermarks will be more imperceptible for human but will induce a large consumption for time and computing resources.
@@
  • Generation-based watermark Previous methods of watermark based on existing texts are based on rule of text [64, 77, 51, 48] which requires complex human design. Zhang et al. [78] propose a novel watermark module without human design which consists of three components: (i) Message encoding module; (ii) Reparameterization module; (iii) Message decoding module. The existing text $T=\{w_{0},w_{1},...,w_{n-1}\}$ and a binary watermark signature $s$ is denoted as input. The message encoding module outputs a watermarked distribution which is $P(T+s)$. $P(T+s)$ will go through the Reparameterization module and the Message decoding module which generates the watermark $s^{*}$. Watermark detection is achieved by comparing $s$ and $s^{*}$. However, our preliminary trial on [78] shows poor watermarking performance with different hyperparameter settings. Therefore, we do not consider this method in our evaluation.
@@
  • Emoji attacks. Kirchenbauer et al. [29] propose an attack that aims to indiscriminate the generative capability of LLMs. We can prompt the LLM to modify its output to some specific form that we desire. Goodside [18] proposes an attack called emoji attack, which can prompt the LLM to generate an emoji after word and then remove these emojis. This attack will randomize the red list for subsequent tokens thus can remove the watermark.
@@
  • Modify attack. For each word in the text, we can duplicate, remove or replace the word to another word with probability $p$.
@@
  • Paraphrase attack. This is an attack that exploits another (paraphrase) language model to rephrase the existing texts. Because of the requirement of accessing another language model, this attack can be expensive to deploy for custom adversaries. Krishana et al. [33] propose a language model which is fine-tuned specially for implementing paraphrase called dipper. In addition, we can also exploit the other models which belong to the Llama family (e.g., WizardLM-2-7B [76] or Llama-3-8B-instruct [4] to implement this attack.
@@
This section provides all the hyperparameters and other experiments setup for all watermark and attack methods.
@@
[…]
For the Lexcial-based watermark, we use the default setting that the similarity threshold is set to 0.5 and the number of candidate synonyms is set to 8 for every word.
@@
The former five attack methods have different experiment setups and we will introduce them respectively. We implement the latter nine attack methods with the same setup. These attack methods will be applied to the model outputs where the watermark has been injected into.
@@
Multi-attack setup. In addition to the initial single attack on the watermarking schemes, we propose novel attack methods that apply multiple attacks to the same watermark. We first apply every two attack methods to the same watermarking schemes and we also use the same attack twice. We then calculate the quality and watermark rate after every combined-attack and select some prominent attacks to apply the multiple attacks. In this work, we screened out the attack methods with the best attack effects.
@@
(a) KGW
(b) Convert
(c) Inverse
(d) Exponential
(e) Unigram
(f) WHITEMARK
(g) UniSpaCh
(h) Linguistic
@@
Robustness score analysis. We first launch individual attacks against watermarking schemes and report the mean values of the robustness scores (Equation 9). As Table 5 shows, we observe that pre-text watermarks are more robust than post-text watermarks in general. For instance, KGW and Exponential reach $0.5220$ and $0.5012$ robustness scores, respectively, while WHITEMARK and UniSpaCh only reach $0.3986$ and $0.4209$ robustness scores. This is expected as the pre-text watermarks usually involve more complex strategies to inject the watermark into the whole text while post-text watermarks only replace a few tokens, which makes them more vulnerable to attacks. However, the robustness score falls significantly short of the benchmark established for the ideal watermark (which should have a 1.0 robustness score), indicating a need for further optimization and improvement.
@@
Quality & Watermark rate. Here we dive deeper into the quality and watermark rate performance of each watermark against different attacks. The results are summarized in Figure 2. Regarding the quality, we find that KGW, Convert, Inverse, Exponential, and UniSpaCh preserve the quality to a large extent after different attacks, e.g., around 0.8. However, for most of the watermarks, the watermark rate drops drastically to less than 0.6, and most of them are near 0.0, which means the watermark cannot be extracted after the attack. This further emphasizes the vulnerability of the watermarks on texts under potential adversarial attacks.
Impact of different attack methods. From now on, focus on KGW to discuss the attacking effectiveness as KGW reaches the best performance in our previous evaluation. We first discover that Typo, Token, Translation and Emoji attacks are more effective than others. For Typo, the attack has a significant impact on the model’s output texts. For Emoji and Token attacks, while these two attacks might only induce minor disruptions in the token generation process, they could potentially impact the distribution of the entire output token list, leading to a significant decrease in the watermark detection of KGW. For Translation and Paraphrase attacks, both forms of attack entirely alter the output text of the model, while preserving the semantic integrity of the said text. Hence, there is a great effect on the output text token which results in a reduction in the watermark rate. Meanwhile, Figure 3(i) demonstrates that the distillation attack is ineffective, as the distilled model retains a watermark rate exceeding 0.65. Moreover, this watermark rate increases with the length of token sequences. This is due to the fact that the distillation approach employed for the model does not inflict any damage on the model’s output.
@@
Figure 4: The quality score of KGW and UniSpaCh watermarking schemes after cobmine-attack.
Figure 5: The watermark rate of KGW and UniSpaCh watermarking schemes after combine-attack.
(a) Attacks with second-level runtime.
(b) Attacks with hour-level runtime.
@@
To more effectively illustrate the vulnerability of machine-generated text watermarks in real-world scenarios, we advocate for the implementation of combined attack strategies. The optimal watermarks for pre-text (KGW) and post-text (UniSpaCh) are chosen as the target watermark algorithm in this experiment and the results are shown in Figure 4 and Figure 5, respectively. From each sub-figure, the rows represent the first attack and the columns represent the second attack.
Results overview. In general, KGW maintains a higher quality and watermark rate than UniSpaCh. For instance, when applying Paraphrase first and Modify later, KGW reaches 0.5726 quality while UniSpaCh only has 0.2542 quality. Meanwhile, it is evident that the columns and rows representing Paraphrase and Translation attacks exhibit superior quality. When Paraphrase and Translation are utilized as the secondary attack, it has been observed that this approach can even lead to an enhancement in the quality of the text. However, the watermark rates are all reduced regardless of the attack order and attack category.
@@
KGW vs UniSpaCh. Through comprehensive analysis of experimental results, we conclude that different watermark schemes exhibit distinct efficacy in withstanding combined attacks. This is due to the inconsistency in the watermark verification methods. For instance, the verification method of KGW calculates a secret key to divide the token’s vocabulary into the green list and the red list and make a statistical analysis of the token percentage that belongs to the two lists. Therefore, Paraphrase attack and translation attacks both disrupt the token statistics by making significant changes to the text, rendering the watermark undetectable. However, both two attacks cannot completely break the specific signals which is a watermark detection mark of UniSpaCh, which induces a slightly higher watermark rate (around 0.01).
@@
Watermark efficiency. We summarize the runtime evaluation results of watermarking schemes in Table 6. We could get the following observations: (1) For the pre-text watermarking schemes, we notice that the watermark injection runtimes are roughly around 0.5 to 1 hour and the detection runtimes are all less than 600 seconds. Taking into account the duration of text generation, the time allocated to watermarking is comparatively minimal. This observation underscores that the operational efficiency of the watermark LLM falls within a tolerable threshold, making it available for real-world scenarios. (2) For the post-text watermarks, we find that there is a great rum-time gap among them. For instance, the experiment results of WHITEMARK and UniSpaCh are both less than 150 seconds. However, the Linguistic watermark is extremely slow with an injection/detection speed of 31.52/56.34 hours. This empirical evidence underscores the significance of the efficiency of injection and detection for watermarking in the overall evaluation of the watermark’s performance. Therefore, we can conclude that even when a watermark exhibits high robustness if it is characterized by low efficiency, its applicability in real-world scenarios is limited. This highlights the need to balance both efficiency and robustness in the watermark design.
@@
Evaluation methodology. As pointed out in [46, 57], an ideal watermarking scheme should exhibit good imperceptibility. In other words, imperceptibility measures how well the watermarking information is hidden such that the watermarking scheme doesn’t affect the readability, coherence, or naturalness of the AI-generated texts. To quantify imperceptibility, we leverage LLM to judge if a text is an AI-generated text that contains watermark and use the classification accuracy as the imperceptibility results. As the texts to be judged are all watermarked texts, consequently, a lower imperceptibility value indicates that LLM cannot recognize the hidden watermark information. Here we regard Llama-3-8B-instruct [4] as the classifier and use the following prompt to classify the texts.
<|begin_of_text|><|start_header_id|>system<|end_header_id|> You are a assistant model to help me to judge if the given text have a watermark and is generated by machine, and please output your judgement ’True’ or ’False’ without any other content after it. ### Text: {text} <|eot_id|><|start_header_id|>user<|end_header_id|> Please provide the judge result. ### Result: <|eot_id|><|start_header_id|>assistant<|end_header_id|>
@@
Another line of works focus on detecting whether the text is real or generated by models by training a classifier or setting threshold for pre-defined metrics [66, 15, 49, 20, 49, 23]. Mitchell et al. [49] develop DetectGPT, which leverages multiple random perturbations to the text and measures the change of the model’s log-likelihood. He et al. [23] propose MGTBench, which is a benchmarking framework including different types of detection methods.

### v2

@@ Abstract
In this paper, to fill this gap, we first systematically comb the mainstream watermarking schemes and removal attacks on machine-generated texts, and then we categorize them into pre-text (before text generation) and post-text (after text generation) classes so that we can conduct diversified analyses. In our experiments, we evaluate eight watermarks (five pre-text, three post-text) and twelve attacks (two pre-text, ten post-text) across 87 scenarios. Evaluation results indicate that (1) KGW and Exponential watermarks offer high text quality and watermark retention but remain vulnerable to most attacks; (2) Post-text attacks are found to be more efficient and practical than pre-text attacks; (3) Pre-text watermarks are generally more imperceptible, as they do not alter text readability, unlike post-text watermarks; (4) Additionally, combined attack methods can significantly increase effectiveness, highlighting the need for more robust watermarking solutions. Our study underscores the vulnerabilities of current techniques and the necessity for developing more resilient schemes.
@@ Introduction
Our work. In this paper, we fill the gap in the literature by systematically categorizing watermarks and attacks into pre-text and post-text classes, where the former involves watermark injection or disruption before text generation, and the latter does so after text generation. We consider eight watermarks (five pre-text and three post-text) and twelve attacks (two pre-text and ten post-text), resulting in a total of 87 possible scenarios (see Table 3). Our evaluation shows that KGW [29] and Exponential [1] are the two best watermarks by providing good text quality and relatively high watermark rates after different attacks compared to other watermarks. For instance, KGW and Exponential achieve over 0.5 robustness score while Linguistic [77] watermarks only reach around 0.3 robustness score (see Table 5 for more details). However, they are still not robust against most of the attacks. For example, KGW only reaches a 0.0349 watermark rate under Paraphrase attacks. This suggests that existing watermarks are still vulnerable to various attacks. Regarding efficiency, both KGW and Exponential are relatively efficient in watermark injection and detection. In terms of attacks, we find that post-text attacks are generally more efficient than pre-text attacks since they do not require modification of the model’s weights, making them more practical threats to current watermarks. Concerning imperceptibility, pre-text watermarks usually perform better as they are hidden in the token distributions and do not affect text readability, whereas post-text watermarks involve adding, removing, or replacing tokens, making them easier to detect.
@@
  • We perform the most comprehensive (thus far) evaluation of the performance of watermarked machine-generated text under adversarial attacks.
@@
Following [53], we model the generation process of LLMs into two steps: probability prediction and token selection. First, given a sequence of tokens $x_{1:n}$ (e.g., the input prompt) where $x_{i}\in\{1,...,|V|\}$ is a token and $|V|$ is the size of vocabulary $V$, LLM calculates the conditional probability of the next token as $Pr(x_{n+1}|x_{1:n})$. Then, LLM uses a specific sampling strategy to select the word from the vocabulary as the final token, e.g., Greedy Sampling confirms the next token by $x_{n+1}=\arg\max_{x}Pr(x|x_{1:n})$. Naturally, LLM generates a $l$ length text through computing $Pr(x_{1:n+l})=\prod_{i=1}^{n+l}Pr(x_{i}|x_{1:i-1})$.
@@
Black-box & White-box. A watermarking scheme is classified as white-box if it relies on the model’s internal parameters during watermark injection. Conversely, a scheme is black-box if it does not access model parameters. Notably, schemes accessing the last hidden state of an LLM are also considered black-box.
@@
First of all, we clarify that we focus on the watermarking schemes which are based on the generated texts in this paper. Compared with the watermarking schemes which need to embed the watermarks by updating the model parameters [56, 36], these text-based schemes can be adapted to any causal language models. Meanwhile, we consider two types of watermarks, i.e., pre-text watermarks and post-text watermarks, where the former injects the watermark before or during text generation and the latter injects the watermark after the text generation.
The methods of pre-text watermarks can be divided into two categories based on the process used for generating text with an LLM [41]. For instance, this includes modifying the logits [29, 79, 40, 35, 45, 24, 43] and modifying the token sampling strategy during the inference phase [34, 1, 9]. The methods of post-text watermark can be divided into four categories based on the granularity of text modification [41], which are format-based watermark [58, 64], lexical-based watermark [77, 51], synaptic-based watermark [70, 48], and generation-based watermark [78, 2]. To comprehensively discuss the robustness of various types of watermarking schemes, we discuss eight LLM watermarking schemes in this paper, which are summarized in Table 1.
@@
Pre-text watermarks can be categorized into token sampling-based and logits modification-based watermarks where the former modifies the token sampling strategy and the latter modifies logits during the inference phase.
@@
  • Convert bit-string (Convert) [9] aims to develop undetectable watermarks. This watermark can be viewed as the hash operation on the generated text. Specifically, they first encode the generated text into a string with binary bits. Then, they use a {0,1} hash function $h$ and sample tokens $x_{j}$ with preference for those satisfying $h(x_{j})=1$. More tokens should hash to 1 than to 0 in watermarked text. The probability of a token being encoded to 1 is defined as $p_{j}(1)$. The watermarked model will output $x_{j}=1$ if $u_{j}\leq p_{j}(1)$ and $x_{j}=0$ otherwise. Here $u_{j}$ is a real number sampled from [0,1]. The probability that the watermarked model output $x_{j}=1$ is $p_{j}(1)$. For detection, they will compute a score $s(x_{j},u_{j})=\ln\frac{1}{u_{j}}$ if $x_{j}=1$ and $s(x_{j},u_{j})=\ln\frac{1}{1-u_{j}}$ otherwise. Given a string $(x_{1},...,x_{L})$, the sum scores of detection is defined as $score=\sum_{j=1}^{L}s(x_{j},u_{j})$. The score of the watermarked text should be significantly higher than the score of the normal text.
@@
  • Inverse token sampling (Inverse) [34] is similar to exponential. This strategy calculates the watermark scores from the next position of the watermark key $\xi$, $\xi=\xi^{(1)},\xi^{(2)},...,\xi^{(m)}$ where $\xi^{(i)}\in[0,1]^{\lvert\mathcal{V}\rvert}$ and m should exceed the max length of generated text, $\mathcal{V}$ is the vocabulary of tokens. Finally, the watermark strategy is defined as:
@@
  • KGW. Kirchenbauer et al. [29] propose a watermarking scheme without retraining the model. Specifically, during the inference process, a secret key can be calculated by the previous token and we can use the secret key to partition the vocabulary into a green list and a red list. Then, we add a bias to each green list logit before the Softmax function to calculate the probability distribution. We can detect the watermark by making a statistical test to the proportion of tokens in the text that belong to the green list. Finally, the statistical measure of the z-score is defined as:
@@
    where $|s|_{G}$ presents the number of green tokens, $L$ presents the total number of tokens in the text and $\gamma$ is a configuration parameter that presents the proportion of green list tokens in the whole vocabulary. We consider the text to be watermarked if its z-score exceeds the pre-defined threshold.
@@
Post-text watermarks constitute a category of schemes wherein watermarks are appended to existing texts. The predominant approach for this type of watermark involves the modification of the existing text to incorporate the watermarking information. We implemented three watermark schemes belonging to this category in this paper, and we divided the three watermarks into two categories which are Format-based watermark and Lexical-based watermark according to the granularity of modifications.
Format-based watermark. The Format-based watermark derives from an image watermark technology [5]. It does not modify the content of text but changes the format of text that are imperceptible for humans, thereby implementing the watermark injection.
  • WHITEMARK [64] exploits the fact that the Unicode of text has several codepoints for whitespace. They can replace the origin whitespace (e.g. U+0020) with another new whitespace (e.g. U+2004) without degrading the quality of the text. The watermark can be detected by calculating the probability of new whitespace.
@@
Lexical-based watermark. This category of watermarks applies word-level modifications by replacing the random sampling words with other candidate synonyms without affecting the structure of the whole sentence. This category of watermarks will be more imperceptible for humans but will induce a large consumption of time and computing resources.
@@
  • Generation-based watermark Previous methods of watermark based on existing texts are based on the rule of text [64, 77, 51, 48] which requires complex human design. Zhang et al. [78] propose a novel watermark module without human design which consists of three components: (i) Message encoding module; (ii) Reparameterization module; (iii) Message decoding module. The existing text $T=\{w_{0},w_{1},...,w_{n-1}\}$ and a binary watermark signature $s$ is denoted as input. The message encoding module outputs a watermarked distribution which is $P(T+s)$. $P(T+s)$ will go through the Reparameterization module and the Message decoding module which generates the watermark $s^{*}$. Watermark detection is achieved by comparing $s$ and $s^{*}$. However, our preliminary trial on [78] shows poor watermarking performance with different hyperparameter settings. Therefore, we do not consider this method in our evaluation.
@@
  • Emoji attacks. Kirchenbauer et al. [29] propose an attack that aims to indiscriminate the generative capability of LLMs. We can prompt the LLM to modify its output to some specific form that we desire. Goodside [18] proposes an attack called emoji attack, which can prompt the LLM to generate an emoji after every word and then remove these emojis. This attack will randomize the red list for subsequent tokens and thus can remove the watermark.
@@
  • Modify attack. For each word in the text, we can duplicate, remove, or replace the word with another word with probability $p$.
@@
  • Paraphrase attack. This is an attack that exploits another (paraphrase) language model to rephrase the existing texts. Because of the requirement of accessing another language model, this attack can be expensive to deploy for custom adversaries. Krishana et al. [33] propose a language model that is fine-tuned specially for implementing paraphrase called dipper. In addition, we can also exploit the other models which belong to the Llama family (e.g., WizardLM-2-7B [76] or Llama-3-8B-instruct [4] to implement this attack.
@@
This section provides all the hyper-parameters and other experiments set up for all watermark and attack methods.
@@
For the Format-based watermark, in WHITEMARK, the replaced whitespace is $U+2004$ and the probability of replacement is set to 0.6. In UniSpaCh watermark, the several specific whitespace we used are { $U+2000$, $U+2001$, $U+2004$, $U+2006$, $U+2007$, $U+2008$, $U+2009$, $U+200A$. The probability of replacement is set to 0.6.
For the Lexical-based watermark, we use the default setting that the similarity threshold is set to 0.5 and the number of candidate synonyms is set to 8 for every word.
@@
The former five attack methods have different experiment setups and we will introduce them respectively. We implement the latter nine attack methods with the same setup. These attack methods will be applied to the model outputs where the watermark has been injected.
@@
Multi-attack setup. In addition to the initial single attack on the watermarking schemes, we propose novel attack methods that apply multiple attacks to the same watermark. We first apply every two attack methods to the same watermarking schemes and we also use the same attack twice. We then calculate the quality and watermark rate after every combined attack and select some prominent attacks to apply the multiple attacks. In this work, we screened out the attack methods with the best attack effects.
@@
(a) KGW
(b) Unigram
(c) Convert
(d) Inverse
(e) Exponential
(f) WHITEMARK
(g) UniSpaCh
(h) Linguistic
@@
Robustness score analysis. We first launch individual attacks against watermarking schemes and report the mean values of the robustness scores (Equation 9). As Table 5 shows, we observe that pre-text watermarks are more robust than post-text watermarks in general. For instance, KGW and Exponential reach $0.5220$ and $0.5012$ robustness scores, respectively, while WHITEMARK and UniSpaCh only reach $0.3986$ and $0.4209$ robustness scores. This is expected as the pre-text watermarks usually involve more complex strategies to inject the watermark into the whole text while post-text watermarks only replace a few tokens, which makes them more vulnerable to attacks. However, the robustness score falls significantly short of the benchmark established for the ideal watermark (which should be above 0.95 robustness score), indicating a need for further optimization and improvement.
@@
Quality & Watermark rate. Here we dive deeper into the quality and watermark rate performance of each watermark against different attacks. The results are summarized in Figure 2. Regarding the quality, we find that KGW, Unigram, Convert, Inverse, Exponential, and UniSpaCh preserve the quality to a large extent after different attacks, e.g., around 0.8. However, for most of the watermarks, the watermark rate drops drastically to less than 0.6, and most of them are near 0.0, which means the watermark cannot be extracted after the attack. This further emphasizes the vulnerability of the watermarks on texts under potential adversarial attacks.
Impact of different attack methods. From now on, focus on KGW to discuss the attacking effectiveness as KGW reaches the best performance in our previous evaluation. We first discover that Typo, Token, Translation, and Emoji attacks are more effective than others. For Typo, the attack has a significant impact on the model’s output texts. For Emoji and Token attacks, while these two attacks might only induce minor disruptions in the token generation process, they could potentially impact the distribution of the entire output token list, leading to a significant decrease in the watermark detection of KGW. For Translation and Paraphrase attacks, both forms of attack entirely alter the output text of the model, while preserving the semantic integrity of the said text. Hence, there is a great effect on the output text token which results in a reduction in the watermark rate. Meanwhile, Figure 3(i) demonstrates that the distillation attack is ineffective, as the distilled model retains a watermark rate exceeding 0.65. Moreover, this watermark rate increases with the length of token sequences. This is due to the fact that the distillation approach employed for the model does not inflict any damage on the model’s output.
@@
(a) Quality score.
(b) Watermark rate.
Figure 4: The quality score and watermark rate of KGW and UniSpaCh watermarking schemes after the combined attack.
@@
To more effectively illustrate the vulnerability of machine-generated text watermarks in real-world scenarios, we advocate for the implementation of combined attack strategies. The optimal watermarks for pre-text (KGW) and post-text (UniSpaCh) are chosen as the target watermark algorithm in this experiment and the results are shown in Figure 4(a) and Figure 4(b), respectively. From each sub-figure, the rows represent the first attack and the columns represent the second attack.
Results overview. In general, KGW maintains a higher quality and watermark rate than UniSpaCh. For instance, when applying Paraphrase first and Modify later, KGW reaches 0.5726 quality while UniSpaCh only has 0.2542 quality. Meanwhile, the columns and rows representing Paraphrase and Translation attacks exhibit superior quality. When Paraphrase and Translation are utilized as the secondary attack, it has been observed that this approach can even lead to an enhancement in the quality of the text. However, the watermark rates are all reduced regardless of the attack order and attack category.
@@
KGW vs UniSpaCh. Through comprehensive analysis of experimental results, we conclude that different watermark schemes exhibit distinct efficacy in withstanding combined attacks. This is due to the inconsistency in the watermark verification methods. For instance, the verification method of KGW calculates a secret key to divide the token’s vocabulary into the green list and the red list and make a statistical analysis of the token percentage that belongs to the two lists. Therefore, Paraphrase attacks and translation attacks both disrupt the token statistics by making significant changes to the text, rendering the watermark undetectable. However, both two attacks cannot completely break the specific signals which is a watermark detection mark of UniSpaCh, which induces a slightly higher watermark rate (around 0.01).
@@
Watermark efficiency. We summarize the runtime evaluation results of watermarking schemes in Table 6. We could get the following observations: (1) For the pre-text watermarking schemes, we notice that the watermark injection runtimes are roughly around 0.5 to 1 hour and the detection runtimes are all less than 600 seconds. Taking into account the duration of text generation, the time allocated to watermarking is comparatively minimal. This observation underscores that the operational efficiency of the watermark LLM falls within a tolerable threshold, making it available for real-world scenarios. (2) For the post-text watermarks, we find that there is a great time gap among them. For instance, the experiment results of WHITEMARK and UniSpaCh are both less than 150 seconds. However, the Linguistic watermark is extremely slow with an injection/detection speed of 31.52/56.34 hours. This empirical evidence underscores the significance of the efficiency of injection and detection for watermarking in the overall evaluation of the watermark’s performance. Therefore, we can conclude that even when a watermark exhibits high robustness if it is characterized by low efficiency, its applicability in real-world scenarios is limited. This highlights the need to balance both efficiency and robustness in the watermark design.
(a) Attacks with second-level runtime.
(b) Attacks with hour-level runtime.
@@
Evaluation methodology. As pointed out in [46, 57], an ideal watermarking scheme should exhibit good imperceptibility. In other words, imperceptibility measures how well the watermarking information is hidden such that the watermarking scheme doesn’t affect the readability, coherence, or naturalness of the AI-generated texts. To quantify imperceptibility, we leverage LLM to judge if a text is an AI-generated text that contains a watermark and uses the classification accuracy as the imperceptibility results. As the texts to be judged are all watermarked texts, consequently, a lower imperceptibility value indicates that LLM cannot recognize the hidden watermark information. Here we regard Llama-3-8B-instruct [4] as the classifier and use the following prompt to classify the texts.
<|begin_of_text|><|start_header_id|>system<|end_header_id|> You are an assistant model to help me judge if the given text has a watermark and is generated by a machine, and please output your judgment ’True’ or ’False’ without any other content after it. ### Text: {text} <|eot_id|><|start_header_id|>user<|end_header_id|> Please provide the judge result. ### Result: <|eot_id|><|start_header_id|>assistant<|end_header_id|>
@@
Another line of works focuses on detecting whether the text is real or generated by models by training a classifier or setting threshold for pre-defined metrics [66, 15, 49, 20, 49, 23]. Mitchell et al. [49] develop DetectGPT, which leverages multiple random perturbations to the text and measures the change of the model’s log-likelihood. He et al. [23] propose MGTBench, which is a benchmarking framework including different types of detection methods.